# Language-Guided World Models [Uncaptioned image] A Model-Based Approach to AI Control

### v1

@@ 6 Related work
World models have a rich history dating back to the 1980s (Werbos, 1987). The base architectures of these models have evolved over time, from feed-forward neural networks (Werbos, 1987), to recurrent neural networks (Schmidhuber, 1990b; a; 1991), and most recently, Transformers (Robine et al., 2023; Micheli et al., 2023). In RL settings, world models serve as the key component of model-based approaches, which use these models for training policies to reduce the amount of interactions with real environments. Model-based RL has been successfully applied to a variety of robotic tasks (Finn & Levine, 2017) and video games (Hafner et al., 2019; 2020; 2023). However, the incorporation of language information into world models has been underexplored. Cowen-Rivers & Naradowsky (2020) propose language-conditioned world models but focus on emergent language rather than human language. Poudel et al. (2023) ground natural language to learn representations for world models. The closest to our work is that of Lin et al. (2023) but the authors do not demonstrate the ability to generalize beyond the training environments.
@@
Language information has been incorporated into various aspects of learning. In instruction following (Bisk et al., 2016; Misra et al., 2018; Anderson et al., 2018; Nguyen & Daumé III, 2019), agents are given descriptions of the desired behaviors and learn to interpret them to perform tasks. Language-based learning (Nguyen et al., 2021; Scheurer et al., 2023) employs verbal descriptions and evaluations of agent behaviors to offer richer learning signals than traditional feedback such as numerical rewards or reference actions. Closest to our work is the line of work that exploits language descriptions of environment dynamics to improve policy learning (Narasimhan et al., 2018; Branavan, 2012; Hanjie et al., 2021; Wu et al., 2023a; Nottingham et al., 2022; Zhong et al., 2020). Rather than using texts to directly improve a policy, our work leverages them to enhance a model of an environment. Recently, several papers propose agents that can read text manuals to play games (Wu et al., 2023a; b). Our work differs from these papers in that we aim to build models that capture exactly the transition function of an environment.
@@ 7 Conclusion
In this work, we introduce Language-Guided World Models, which can be adapted through natural language. These models have numerous advantages over traditional observational world models. We demonstrate the challenges of building these models even in a grid-world environment and present an approach that outperforms strong baselines. Our model is still lacking in performance and the grid-world environments we experiment with severely underrepresent the real world. Nevertheless, we hope that this work helps envision the potential of LWMs in enhancing the safety and utility of artificial agents for humans, and inspires future efforts to further develop these models.
@@ Acknowledgements
Appendix A Dataset
Train 1,536 986 101,376 92
NewCombo 896 598 4,480 89
NewAttr 204 319 1,020 88
Dev NewAll 856 1,028 4,280 86
NewCombo 896 587 4,480 90
NewAttr 204 306 1,020 93
Test NewAll 856 1,016 4,280 88
Appendix B ChatGPT parsing prompt
The following is the prompt that we compose for parsing descriptions. We use the “May 3, 2023” release of ChatGPT. We feed to the model one description at a time instead of a whole manual of three descriptions. We ask it to also extract the role and movement pattern, but use only the parsed identity in the GPTHard model. The “ChatGPT identity-parsing” column in Table 4 shows the fraction of games in each split in which ChatGPT correctly identifies all three identities in a game.
@@ Appendix C Training details
Number of decoder token blocks 33 (max. trajectory length + 1)
@@
def init_weights(module):
if isinstance(module, (nn.Linear, nn.Embedding)):
module.weight.data.normal_(mean=0.0, std=0.02)
if isinstance(module, nn.Linear) and module.bias is not None:
%**** 8-appendix.tex Line 150 ****module.bias.data.zero_()
elif isinstance(module, nn.LayerNorm):
module.bias.data.zero_()
module.weight.data.fill_(1.0)
which is evoked by calling self.apply(init_weights) in the model’s constructor. We initialize all models with this scheme, but only EMMA-LWM and OracleParse perform well consistently on various random seeds.
@@ Appendix D Imitation learning experiments
For the online IL setting, we use the win the game rule-based policy (Appendix A) as the expert. For the filtered BC setting, we train an EMMA policy to overfit the test environment. We then use a fully converged checkpoint of the policy as the near-optimal expert, and a not fully converged checkpoint as the suboptimal expert. The former is trained for 10,000 iterations and the latter is trained for 2,000 iterations.
@@
Appendix E Mathematical interpretation of the cross entropy loss
Let $P_{\pi}(h\mid E)$ be the distribution over histories in environment $E$ induced by executing a behavioral policy $\pi$. We evaluate an LWM on environments drawn from a distribution $P_{{\mathcal{S}},{\mathcal{A}}}^{\textrm{eval}}(E)$, which may not necessarily be the same as the training distribution $P_{{\mathcal{S}},{\mathcal{A}}}^{\textrm{train}}(E)$. Let ${\mathcal{B}}_{\textrm{eval}}$ be a set of trajectories sampled using $P_{{\mathcal{S}},{\mathcal{A}}}^{\textrm{eval}}(E)$ and $\pi$. We report $J(\theta;\mathcal{B}_{\textrm{eval}})$ as an estimate of the expected KL divergence between the evaluated and true models under the joint distribution $P(h,E\mid\pi)\triangleq P_{\pi}(h\mid E)P_{{\mathcal{S}},{\mathcal{A}}}^{\textrm{eval}}(E)$:
$J(\theta;\mathcal{B}_{\textrm{eval}})$ $\approx\underbrace{\mathbb{E}_{E\sim P_{{\mathcal{S}},{\mathcal{A}}}^{\textrm{eval}}(\cdot),h\sim P_{\pi}(\cdot\mid E)}[\mathcal{H}(M,M_{\theta}\mid h,\ell_{E})]}_{\text{expected conditional cross entropy between $M$ and $M_{\theta}$}}$
$=\underbrace{\mathbb{E}_{E\sim P_{{\mathcal{S}},{\mathcal{A}}}^{\textrm{eval}}(\cdot),h\sim P_{\pi}(\cdot\mid E)}[\text{KL}(M\mid\mid M_{\theta};h,\ell_{E})]}_{\text{expected conditional KL divergence between $M$ and $M_{\theta}$}}+\underbrace{\mathbb{E}_{E\sim P_{{\mathcal{S}},{\mathcal{A}}}^{\textrm{eval}}(\cdot),h\sim P_{\pi}(\cdot\mid E)}[\mathcal{H}(M\mid h,\ell_{E})]}_{\text{expected conditional entropy of $M$ (constant w.r.t. $\theta$)}}$
Note that the minimum value of this metric is not zero when the environment is stochastic.
Appendix F Extended imaginary imitation learning results (Table 6)

### v2

@@ 6 Related work
World models have a rich history dating back to the 1980s (Werbos, 1987). The base architecture has evolved from feed-forward neural networks (Werbos, 1987), to recurrent neural networks (Schmidhuber, 1990a, b, 1991), and most recently, Transformers (Robine et al., 2023; Micheli et al., 2023). In RL settings, world models are the key component of model-based approaches, which train policies in simulation to reduce the amount of interactions with real environments. Model-based RL has been successful in a variety of robotic tasks (Finn and Levine, 2017) and video games (Hafner et al., 2019, 2020, 2023). However, the incorporation of language information into world models has been underexplored. Cowen-Rivers and Naradowsky (2020) propose language-conditioned world models but focus on emergent language rather than human language. Poudel et al. (2023) incorporate features language into the representations of the model. These approaches, however, do not use language to control a world model.
@@
Language information has been incorporated into various aspects of learning. In instruction following (Bisk et al., 2016; Misra et al., 2018; Anderson et al., 2018; Nguyen and Daumé III, 2019), agents are given descriptions of the desired behaviors and learn to interpret them to perform tasks. Language-based learning (Nguyen et al., 2021; Scheurer et al., 2023) employs language-based feedback to train models. Another line of work uses language descriptions of environment dynamics to improve policy learning (Narasimhan et al., 2018; Branavan, 2012; Hanjie et al., 2021; Wu et al., 2023a; Nottingham et al., 2022; Zhong et al., 2020). Rather than using texts to directly improve a policy, our work leverages them to enhance a model of an environment. Recently, several papers propose agents that can read text manuals to play games (Wu et al., 2023a, b). Our work differs from these papers in that we aim to build models that capture exactly the transition function of an environment.
Compositional generalization for language-guided world models.
Lin et al. (2024) model a variety of text-augmented environments but do not demonstrate the generalizability of their approach in Messenger. Recent work (Zhao et al., 2022; Du et al., 2024; Zhou et al., 2024; Zhang et al., 2024) has developed LWMs with compositional generalizability. While these papers operate on more visually realistic domains than ours, the language they study is simpler, focusing on concepts that correspond to straightforward mappings from input to output such as colors and objects. In contrast, the concepts in Messenger are more intricate, regarding interactions among multiple entities.
@@ 7 Conclusion
We introduce Language-Guided World Models, which can be adapted through natural language. We outline numerous advantages of these models over traditional observational world models. Our model is still lacking in performance and the grid-world environments we experiment with severely underrepresent the real world. Nevertheless, we hope that this work helps envision the potential of LWMs in enhancing the controllability of artificial agents and inspires future efforts to address the compositional generalization challenge.
@@ Acknowledgements
Appendix A GPTHard model
This approach leverages the language-understanding capabilities of ChatGPT. Through few-shot prompting, we instruct this model to determine the identity of the entity mentioned in each manual description. In this approach, we generate only the set of values $\bm{m}^{\textrm{val}}$ as in Eq 1. Instead of learning soft attention, we directly route the values to the identity embeddings. Concretely, the feature vector added to $\bm{i}^{c}_{t}$ in Eq 3 is $\bm{z}^{c}_{t}=\bm{m}^{\textrm{val}}_{j_{c}}$ where $j_{c}$ is the index of the description that mentions entity $c$ according to ChatGPT.
We compose the following prompt for parsing descriptions. We use the “May 3, 2023” release of ChatGPT. We feed to the model one description at a time instead of a whole manual of three descriptions. We ask it to also extract the role and movement pattern, but use only the parsed identity in the GPTHard model. The “ChatGPT identity-parsing” column in Table 4 shows the fraction of games in each split in which ChatGPT correctly identifies all three identities in a game. Note that the OracleParse model uses the ground-truth parses rather than these parses.
@@ Appendix A GPTHard model
Appendix B Dataset
Train 1,536 986 101,376 92
NewCombo 896 598 4,480 89
NewAttr 204 319 1,020 88
Dev NewAll 856 1,028 4,280 86
NewCombo 896 587 4,480 90
NewAttr 204 306 1,020 93
Test NewAll 856 1,016 4,280 88
@@ Appendix C Training details
Number of decoder token blocks 33
@@
{minted}
python def init_weights(module): if isinstance(module, (nn.Linear, nn.Embedding)): module.weight.data.normal_(mean=0.0, std=0.02) if isinstance(module, nn.Linear) and module.bias is not None: module.bias.data.zero_() elif isinstance(module, nn.LayerNorm): module.bias.data.zero_() module.weight.data.fill_(1.0) which is evoked by calling self.apply(init_weights) in the model’s constructor. We initialize all models with this scheme, but only EMMA-LWM and OracleParse perform well consistently on various random seeds.
Compute resources.
Experiments were primarily run on a cluster of NVIDIA RTX2080 GPUs, and each experiment was run on a single device. To generate Table 1, we trained each world model for 24 GPU hours, 5 seeds each. To generate Table 3 and 6, we trained each of the 5 world models on each of the 90 games (3 difficulties for 30 game configurations) using the 3 different downstream policy training strategies, with each game being 12 GPU hours.
@@ Appendix D Imitation learning experiments
For the online IL setting, we use the win the game rule-based policy (Appendix B) as the expert. For the filtered BC setting, we train an EMMA policy to overfit the test environment. We then use a fully converged checkpoint of the policy as the near-optimal expert, and a not fully converged checkpoint as the suboptimal expert. The former is trained for 10,000 iterations and the latter is trained for 2,000 iterations.
@@
Appendix E Extended results
Table 6 presents the results of all the models in the simulation of plan discussion (§ 5.3).